A business supplies the following figures about its activities.

Fixed                                        $300,000
Variable Cost                         $ 20 per unit
Forecast output (sales)      20,000 Units
Selling Price                           $ 50 per unit

Required
Illustrate by means of breakeven chart

I. The breakeven point
ii. The profit at full capacity
iii. The margin of safety
Answer #1

Data table to be made as under:

No. of units Fixed cost Total Variable cost Total cost (Fixed + Variable) Total Revenue
                  -   $        300,000 $                   -   $        300,000 $                       -  
           5,000 $        300,000 $        100,000 $        400,000 $            250,000
         10,000 $        300,000 $        200,000 $        500,000 $            500,000
         15,000 $        300,000 $        300,000 $        600,000 $            750,000
         20,000 $        300,000 $        400,000 $        700,000 $        1,000,000
         25,000 $        300,000 $        500,000 $        800,000 $        1,250,000
         30,000 $        300,000 $        600,000 $        900,000 $        1,500,000

Breakeven chart is pasted below:

$1,600,000 $1,400,000 $1,200,000 $1,000,000 S800,000 S600,000 $400,000 $200,000 S- 10,000 30,000 5,000 15,000 20,000 25,000 N

i) The breakeven point

breakeven point in units = Fixed cost / Contribution margin per unit

breakeven point in sales = Fixed cost / Contribution margin ratio

Contribution margin per unit = Sales price per unit - Variable cost per unit

=$ 50 - 20 = $ 30

Contribution margin ratio = Contribution margin per unit / Sales price per unit * 100

= 30 / 50 * 100 = 60%

breakeven point in units = $ 300,000 / 30 = 10,000 units

breakeven point in sales = $ 300,000 / 60% = $ 500,000

ii) The profit at full capacity is $ 300,000

Calculations to be made as under:

Per unit For 20,000 units
Sales                  50             1,000,000
Variable cost                  20                 400,000
Contribution margin                  30                 600,000
Fixed cost                 300,000
Profit                 300,000

iii) The margin of safety

Margin of safety = (Current sales level - Breakeven sales) / Current sales level * 100

= ($ 1,000,000 - 500,000) / 1,000,000 * 100

= 50%

Margin of safety = 50%